1.For a body dropped from a height, h, the speed, V,with which it strikes the ground depends only on h

and g and a dimensionless numerical factor. Obtain an expression connecting v, h and g.


1
Expert's answer
2021-10-28T13:14:18-0400

"a=g"

"v=gt+v_0"


if "v_0=0" :

"v=gt"


"h_t=h-gt^2\/2"

where h is height at t=0,

ht is height at moment t


when body strikes the ground:

"h_t=0"

"t=\sqrt{\frac{2h}{g}}"


speed, when body strikes the ground:


"v=g\sqrt{\frac{2h}{g}}=\sqrt{2gh}"